ShoaibChaudhry:

Solve log4 (y − 9) + log4(3) = log4(81).

supie:

@jhonyy9

jhonyy9:

use log(a)b +log(a)c = log(a)b×c log(a)b -log(a)c = log(a)b/c

jhonyy9:

rewrite 81 like power of 3

jhonyy9:

suppose log4(y-9) is the first term